OpenStudy (anonymous):

find the real solutions 4(x+1)^2+23(x+1)+15=0

OpenStudy (agent0smith):

You can factor it as is, or... Make a substitution to make this easier. Let u = x+1 then your equation is 4u^2 + 23u + 15 = 0 try factoring this.

OpenStudy (anonymous):

u(4u+23)+-15 , U+ -15 or 2

OpenStudy (anonymous):

mistake U= -15 or 2

OpenStudy (agent0smith):

That doesn't look factored right at all. You need to first multiply the first number by the last number, 4u^2 + 23u + 15 = 0 4*15=60. Now find two numbers which multiply to make 60, and add up to the middle number, 23.

OpenStudy (anonymous):

20 and 3

OpenStudy (anonymous):

so now you have (4u+3)(u+5)

OpenStudy (agent0smith):

(4u+3)(u+5) = 0 now just replace u with x+1 and solve for x.

OpenStudy (anonymous):

thanks, I got the answer

OpenStudy (anonymous):

ttheres one more problem........solve by factoring 4x-11=3/x

OpenStudy (anonymous):

arent i suppose to get rid of the x in the bottom?

OpenStudy (agent0smith):

Yes, multiply both sides by x. Then subtract 3 on both sides, then factor it.

OpenStudy (agent0smith):

You'll be factoring 4x^2 - 11x - 3 = 0

OpenStudy (anonymous):

So -1/4 and 3...your awesome, I truly appreciate it. I understand now